A property has a market value of $750,000. The taxing authority taxes 70% of the value and the property has a homestead exemption of $25,000. The county mill rate is 8.4 and the town mill rate is 11.2. If the taxes are $15,000, what is the board of education’s mill rate?
Answer #1

Mill Rate = Tax / (Property value * Taxable Value - Exemption) *100% = 15000 / (750000*70%-25000) *100% = 3%